Ratios are mostly calculated using data drawn from the financial statements of a firm. However, another group of ratios, called market-based ratios, relate to a firms observable market value, stock prices, and book values integrating information from both the market and the firms financial statements. Consider the case of Green Caterpillar Garden Supplies Inc.: Green Caterpillar Garden Supplies Inc. just reported earnings after tax (also called net income) of $95,000,000, and a current stock price of $17.50 per share. The company is forecasting an increase of 25% for its after-tax income next year, but it also expects it will have to issue 2,800,000 new shares of stock (raising its shares outstanding from 5,500,000 to 8,300,000).

If Green Caterpillars forecast turns out to be correct and its price-to-earnings (P/E) ratio does not change, what does the companys management expect its stock price to be one year from now? (Round any P/E ratio calculation to four decimal places) $14.50 per share O $17.50 per share $10.88 per share O $18.13 per share One year later, Green Caterpillars shares are trading at $52.08 per share, and the company reports the value of its total common equity as $24,070,000. Given this information, Green Caterpillars market-to-book (M/B) ratio is Is it possible for a company to exhibit a negative EPS and thus a negative P/E ratio? O Yes O No Which of the following statements is true about market value ratios? O Companies with high research and development (R&D) expenses tend to have low P/E ratios. Companies with high research and development (R&D) expenses tend to have high P/E ratios
Answer #1

1.Original P/E ratio = Price Per share/(Earnings Per Price) = 17.50/(95,000,000/5,500,000) = 1.0132

New Price Per Share/(New Earnings Per Share) = 1.0132
New Price Per Share = 1.0132*(95,000,000*1.25)/(8,300,000) = 14.50

Option a is correct.

2. M/B ratio = 52.08/(24,070,000/830,000) = 17.96

3. Yes

4.Option b . High P/E ratio

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well